what is th nth term and amt of 3,12,27,48
Answer by greenestamps(13200)



The given numbers are...
3 = 3(1^2)
12 = 3(2^2)
27 = 3(3^2)
48 = 3(4^2)

So the n-th term is PROBABLY 3(n^2).

However, the problem as stated is invalid, because ANY following numbers would make a valid sequence.